OpenStudy (anonymous):

How much heat, in kilojoules, is released from burning all of the methane from one tank of LNG?When 1.0 mole of methane gas burns with oxygen to produce carbon dioxide and water vapor, 883 kJ of heat is produced. At STP, methane gas has a density of 0.715 g/L. For transport, the natural gas is cooled to -163 ∘C to form liquefied natural gas (LNG) with a density of 0.45 g/mL. A tank on a ship can hold 6.8 million gallons of LNG.

OpenStudy (shiraz14):

1 US gal = 3785.41mL 6.8 million US gal = 2.57408 * 10^10 mL Mass of LNG present = 2.57408 * 0.45 * 10^10 g = 1.158336 * 10^10 g Volume of methane at STP = (1.158336 * 10^10) / (0.715) L At STP, 1 mole of gas occupies a volume of 22.4L. Hence, we have: Number of moles of methane = (1.158336 * 10^10) / (22.4 * 0.715) mol Amount of heat energy released from burning 1 tank of LNG = [(1.158336 * 10^10 * 883) / (22.4 * 0.715)] kJ = 6.386*10^11 kJ (Answer)